OpenStudy (anonymous):

Which expression is equivalent to (9^{-2})^{8}?

OpenStudy (anonymous):

Which expression is equivalent to \[(9^{-2})^{8}\]?

OpenStudy (denonakavro):

Hey lets start with the parenthesis \[9^{-2} = \frac{ 1 }{ 81 }\]

OpenStudy (mathmale):

Agreed. Work shown within parentheses has to be done first. Using 1/81 instead of 9^(-2), re-write (9^{-2})^{8}.

OpenStudy (mathmale):

Kristin: We'd like to have your input.

OpenStudy (anonymous):

Ok so \[\frac{ 1 }{ 81 } * \frac{ 8 }{ 1 } = \frac{ 8 }{ 81 }\] Is that right? :/

OpenStudy (danica518):

its to the power of 8

OpenStudy (danica518):